Conference Site Proposal

Draft of September 23, 1994 Submitted by:

Site:

Contact Person: Year:

Telephone: Cost Commitment:

(indicate currency used)

\underline{Provisional Organizing Committee Members} Chair

Member  Member  Member
Member  Member  Member
Member  Member  Member

Hotel/Conference Location Yes No Comments/Price

At a university?  □  □
Are there hotels nearby?  □  □
At a hotel?  □  □
If so, are other hotels nearby?  □  □

Deadline for signing contract?

Are hotel rates available now?  □  □

What types of restaurants are nearby?

Airport Recommended

Distance to the conference site?  □  □

What type of airport transportation is available:

Shuttle service  □  □
Limousine  □  □
Taxi  □  □
Rental Car  □  □

Conference-Specific Requirements Yes No Comments/Price

How many meeting rooms are available What are the meeting room sizes?

Are computer and/or training facilities locally available?  □  □
Is conference staff provided by the hotel/university staff?  □  □

Are the following offered by the conference host or located nearby:

Electronic mail access  □  □
Fax services  □  □
Photocopying services  □  □

Does the conference site provide services for:

Wheelchair access  □  □
Sign language interpreters  □  □
Foreign language interpreters  □  □

Miscellaneous Yes No Comments/Price

Are maps available of interesting sites to see?  □  □
Is a Chamber of Commerce tourist package available?  □  □

Return this Conference Site Planning Checklist (short form) to tugcpc-l@irlearn.ucd.ie